An investor with $10,000 wishes to invest part of it at 11% simple interest and the rest at 12% simple interest. How much should she invest at 11% if she is to receive a total of $1115 in interest from boht investments that year?
Thanks!

Answer by checkley71(8403)

.11x+.12(10,000-x)=1,115 or .11x+1,200-.12x=1,115 or -.01x=-85 or x=85/.01 or
x=8,500 invested @ 11% & 10,000-8,500=1,500 @ 12%.
proof .11*8,500+.12*1,500=1,115 or 935+180=1,115 or 1,115=1,115